OpenStudy (anonymous):

What if a line passes through the origin and it has to be parallel to x + y = 10

OpenStudy (anonymous):

What can you say about the slope of two parallel lines?

OpenStudy (anonymous):

That they are the same

OpenStudy (anonymous):

Ok , can you determine the slope of x + y = 10?

OpenStudy (anonymous):

No I cant

OpenStudy (anonymous):

Slope-Intercept Form is y=mx+b where m is the slope sample 2x + y = 3 you will solve for y in terms of x, so this will become y = -2x + 3 so the slope in this example is -2. can you do it now?

OpenStudy (anonymous):

K

OpenStudy (anonymous):

Is it -y = x + 10?

OpenStudy (anonymous):

check it again .. x + y = 10 ----- >

OpenStudy (luigi0210):

Once you find the slope of that equation, everything else is pretty simple. It is asking for an equation through the origin, meaning (0, 0) So use: \[y-y_{1}=m(x-x_{1})\]

OpenStudy (luigi0210):

In slope-intercept form, the y can not be negative

OpenStudy (anonymous):

Oh k

OpenStudy (luigi0210):

So the equation you had was okay, it just needed to be multiplied by a -1.

OpenStudy (anonymous):

No it's not, its should be -10 if that's the case.

OpenStudy (anonymous):

Y?

OpenStudy (luigi0210):

What? -10?

OpenStudy (anonymous):

cos you said his equation was ok, just multiply by -1, i don't think so

OpenStudy (luigi0210):

Oh, I see what you mean. Yea, show us your final answer @hubbabubba for the equation

OpenStudy (anonymous):

You guys confused me sooo bad omg I quit